# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

### v1

@@ Abstract
We present a detailed investigation of photometric, spectroscopic, and polarimetric observations of the Type II SN 2023ixf. The early detection of highly-ionized “flash” features, rapid ascent in ultraviolet flux coupled with the blueward shift in near-ultraviolet colors and temperature provides compelling evidence for a delayed shock breakout from a confined dense circumstellar matter (CSM) enveloping the progenitor star. The temporal evolution of polarization in the SN 2023ixf phase revealed three distinct peaks in polarization evolution at 1.4 d, 6.4 d, and 79.2 d, indicating an asymmetric dense CSM, an aspherical shock front and clumpiness in the low-density extended CSM, and an aspherical inner ejecta/He-core. SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry in the early and late evolution. The argument for an aspherical shock front is further strengthened by the presence of a high-velocity broad absorption feature in the blue wing of the Balmer features in addition to the P-Cygni absorption post 16 d. Hydrodynamical light curve modeling indicated a progenitor mass of 10 with a radius of 470 and explosion energy of , along with 0.06 of . The modeling also indicated a two-zone CSM: a confined dense CSM extending up to , with a mass-loss rate of 10-2 and an extended CSM spanning from to with a mass-loss rate of 10-4 . The early nebular phase observations display an axisymmetric line profile of [O I], and red-ward attenuation of the emission of H post 125 days, marking the onset of dust formation.
@@ 1 Introduction
Supernovae (SNe) are catastrophic explosions that are crucial in shaping the evolution of the Universe through chemical enrichment and dust formation. Massive stars ( 8 , Heger et al., 2003) reach the termination of their evolutionary phase upon exhaustion of their nuclear fuel and undergo core-collapse supernovae (CCSNe). SNe II form a subset of CCSNe that show Balmer features in their spectral sequence (Filippenko, 1997). The progenitors of these SNe are confirmed to be red supergiants (RSGs) through pre-explosion imaging of several SNe II in nearby galaxies (Smartt, 2009; Van Dyk, 2017).
@@
Massive stars are thought to have steady stellar winds driven by radiation pressure (Chiosi & Maeder, 1986) with the Galactic RSGs having an observed mass-loss rate of (Beasor et al., 2020) in agreement with such winds. However, early spectral observations of SNe II have shown a short-lived ‘flash’ phase with narrow high-ionization features. Such features were first studied extensively in SN 1998S (Gerardy et al., 2000) as they trace the confined CSM arising from late stages of mass-loss in RSGs (Yaron et al., 2017), indicating a much higher mass loss rate than steady winds (Dessart et al., 2022). A significant fraction of SNe II discovered by Zwicky Transient Facility (ZTF) exhibit such spectroscopic signatures of early interaction during their evolution ( 36%, Bruch et al., 2023).
@@
The effects of a dense CSM driven by high mass loss rates have been known to delay shock breakout (Förster et al., 2018). It occurs when the shock encounters a medium of significantly low optical depth, allowing the radiation to escape in front of the shock (Waxman & Katz, 2017). The breakout is the first electromagnetic signature in CCSNe and happens within a few hours after core-collapse for a typical RSG (Matzner & McKee, 1999). However, if the massive star is engulfed in a compact, dense CSM, the shock breakout leads to an elongated and brighter electromagnetic signature (Ofek et al., 2010; Chevalier & Irwin, 2011). The emergence of high-cadence all-sky surveys (Förster et al., 2018; Subrayan et al., 2023) coupled with prompt follow-up observations has photometrically revealed that fast-rising light curves are primarily reproduced via interaction with CSM. Such an interaction with CSM complicates the classification of SNe II (Singh et al., 2019a). SNe II are primarily classified into two main sub-types (Barbon et al., 1979) based on their light curve decline rates: a Type IIP SN with a slow decline and a Type IIL SN with a steep linear decline during the plateau phase. The primary distinction between Type IIP/L SNe arose from the divide between the outer hydrogen envelope mass. However, studies by Moriya et al. (2017, 2018); Morozova et al. (2017) have suggested that a higher decline rate during the plateau phase is primarily driven by interaction with CSM, its compactness, and geometry (Andrews et al., 2019; Jacobson-Galán et al., 2022). Chugai (2021) also indicated that higher CSM mass and wind density have been inferred for steeper declining SNe IIL compared to SNe IIP.
@@
Detailed studies of highly interacting SNe II such as SN 2013fs (Yaron et al., 2017), SN 2020pni (Terreran et al., 2022), SN 2020tlf (Jacobson-Galán et al., 2022) etc, have suggested a much higher mass-loss rate for their RSG progenitors, i.e., . Since steady stellar winds cannot propel such a high mass loss rate in an RSG (Puls et al., 2008), the paradigm has shifted towards a more complex scenario in recent years. These mechanisms involve eruptive bursts of mass loss (Smith & Owocki, 2006), extensive RSG mass loss through surface instability (Yoon & Cantiello, 2010), wave-driven mass loss during the end stages of nuclear burning (Quataert & Shiode, 2012) and/or turbulent convection driven by dynamical instability (Smith, 2014). The complex mass-loss mechanisms markedly influence the properties of a CCSN through the diverse pathways of forming CSM and its resulting geometrical configuration.
@@
This work presents the multi-wavelength analysis of SN 2023ixf through photometric, spectroscopic, and polarimetric monitoring. SN 2023ixf was discovered by Koichi Itagaki (Itagaki, 2023) at JD 2460084.4 and reported on Transient Name Server (TNS) followed by a prompt classification as a Type II SN by Perley & Gal-Yam (2023). The data spanning the first 20 days of our study have already been published in Teja et al. (2023a) discussing the early photometric and spectroscopic analysis. We adopt the explosion epoch as estimated by Teja et al. (2023a), i.e., JD 2460083.315, and the epoch described hereafter in the text is described with respect to this date.
We provide an overview of the host galaxy’s properties in Section 2, and summarize the details on observational data acquisition and reduction in Section 3. The photometric properties of SN2023ixf are analyzed and discussed in Section 4, followed by a discussion on its spectroscopic evolution in Section 5. Section 6 delves into the temporal evolution of the polarimetric characteristics. Section 7 describes the hydrodynamical light curve modeling to infer the properties of the progenitor and its multi-faceted CSM. We qualitatively discuss the presence of a confined CSM and the signatures of asphericity in the CSM/ejecta in order to understand the progenitor of SN 2023ixf in Section 8 and summarize our results in Section 9.
@@ 2 Host Galaxy - M101
SN 2023ixf exploded in the outer spiral arm of the face-on spiral Galaxy, M 101 (NGC 5457), located in the constellation Ursa Major at a redshift of z = 0.0008046 obtained from NASA Extragalactic Database (NED111https://ned.ipac.caltech.edu). SN 2023ixf is the 2nd nearest CCSN in this millennium after SN 2004dj. SN 2023ixf lies proximal to one of the giant H II regions in M 101, namely NGC 5461 (see Figure 1) and in the immediate neighbourhood of the H II region #1086. Using emission-line diagnostics of the spectra of H II region #1086, (Van Dyk et al., 2023) estimated an oxygen abundance through of 8.43 12 + log[O/H] 8.86, which equates to a metallicity of 0.10 Z 0.020 close to the site of the SN. We adopt a mean distance of 6.82 0.14 Mpc ( = 29.17 0.04 mag), a Galactic reddening of = 0.00770.0002 mag, the host reddening of = 0.031 0.011 mag, and the total reddening of = 0.039 0.011 mag for SN 2023ixf from Teja et al. (2023a). These values are consistent with the values estimated by Van Dyk et al. (2023); Smith et al. (2023).
@@ 3.1 Ultraviolet/Optical/Near-Infrared Photometry
The optical photometric observations of SN 2023ixf were performed using the robotic 0.7-m GROWTH-India telescope (GIT, Kumar et al., 2022) located at the Indian Astronomical Observatory (IAO) in Hanle, India. The observations were carried out in the Sloan Digital Sky Survey (SDSS) filters beginning 2023 May 20 UT. The data were processed with the standard GIT image processing pipeline described in Kumar et al. (2022), and the steps followed are described in Teja et al. (2023a). We also carried out photometric observations in using a 0.36-m Schmidt Cassegrain telescope (Celestron EdgeHD 1400) at the Home observatory in Nayoro, Hokkaido, utilizing a CCD FLI ML1001E camera with an IDAS filter (standard system). The data reduction and aperture photometry were carried out using the software MIRA Pro 64. The calibration was performed using the stars from APASS catalog (Zacharias et al., 2013). Additionally, optical photometric observations in bands were also carried out using Atik 460 EX Mono CCD mounted on the 0.61-m Vasistha telescope at Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), Sitapur, ICSP, Kolkata. Photometric calibrations were done using Tycho software and the ATLAS catalog (Tonry et al., 2018). -band imaging was also carried out using the 0.51-m telescope at Oku Observatory, Okayama, using the SBIG Camera STXL-6303. The data reduction was performed according to the standard procedure using IRAF.
@@
The Ultraviolet Optical Telescope (UVOT; Roming et al., 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) monitored SN 2023ixf extensively beginning May 21, 2023. We reduced the publicly available data obtained from Swift Archives444Swift Archive Download Portal. Photometry was performed using the UVOT data analysis software in HEASoft, following the procedure described in Teja et al. (2022). Upon using a 5 aperture, a significant part of the nearby H II contaminates the flux, especially in the ultraviolet (UV) bands at later epochs ( 20 d). We used the archival data for the host M101 obtained under OBSID 00035892001, available in the Swift archive, as template images for removing the host contribution. The flux obtained at the SN site in the template images is comparable to the late-phase fluxes in all the swift bands. We employed Swift_host_subtraction555https://github.com/gterreran/Swift_host_subtraction code (Brown et al., 2009, 2014a) to remove the host contribution.
@@ 3.2 Optical Spectroscopy
Low-resolution optical spectroscopic observations of SN 2023ixf were carried out using the Himalayan Faint Object Spectrograph (HFOSC) instrument mounted on the 2-m Himalayan Chandra Telescope (HCT) at IAO (Prabhu, 2014). The HFOSC observations were performed in grisms Gr#7 (3500-7800 Å, R 500) and Gr#8 (5200-9250 Å, R 800). The spectroscopic data were reduced in a standard manner using the packages and tasks in IRAF with the aid of the Python scripts hosted at RedPipe (Singh, 2021). Optical spectroscopic observations were also performed using the Kyoto Okayama Optical Low-dispersion Spectrograph with optical-fiber Integral Field Unit (KOOLS-IFU, Matsubayashi et al., 2019) mounted at the 3.8-m Seimei Telescope (Kurita et al., 2020) located in Okayama Observatory, Kyoto University, Japan. The KOOLS-IFU observations were carried out using VPH-blue (4100-8900 Å, R 500). The data reduction was performed using the Hydra package in IRAF and a reduction software developed for KOOLS-IFU data666http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools. Arc lamps of Hg, Ne, and Xe were used for wavelength calibration. We obtained additional optical spectroscopic data using the TriColor CMOS Camera and Spectrograph (TriCCS) installed on the 3.8-m Seimei telescope having a wavelength coverage of 4000–10500 Å (R 700). The spectroscopic observations were also carried out using the spectrograph installed on the 0.4-m reflector at the Fujii Kurosaki Observatory (FKO) in Okayama, Japan, with a resolution of and a wavelength coverage of 4000-7800 Å. The spectroscopic data were reduced using standard IRAF tasks. We used the L. A. Cosmic pipeline (van Dokkum, 2001) to remove cosmic rays. We used arc lamps (Hg, Ne, and Xe) for wavelength calibration, and the flux was calibrated using spectrophotometric standards. All the spectra have been continuum calibrated with respect to photometry and corrected for the redshift of the host galaxy.

### v2

@@ Abstract
We present a detailed investigation of photometric, spectroscopic, and polarimetric observations of the Type II SN 2023ixf. Earlier studies have provided compelling evidence for a delayed shock breakout from a confined dense circumstellar matter (CSM) enveloping the progenitor star. The temporal evolution of polarization in SN 2023ixf revealed three distinct peaks in polarization evolution at 1.4 d, 6.4 d, and 79.2 d, indicating an asymmetric dense CSM, an aspherical shock front and clumpiness in the low-density extended CSM, and an aspherical inner ejecta/He-core. SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry in the early and late evolution. The argument for an aspherical shock front is further strengthened by the presence of a high-velocity broad absorption feature in the blue wing of the Balmer features in addition to the P-Cygni absorption post 16 d. Hydrodynamical light curve modeling indicated a progenitor mass of 10 with a radius of 470 and explosion energy of , along with 0.06 of , though these properties are not unique due to modeling degeneracies. The modeling also indicated a two-zone CSM: a confined dense CSM extending up to , with a mass-loss rate of 10-2 and an extended CSM spanning from to at least with a mass-loss rate of 10-4 , both assuming a wind-velocity of 10 . The early nebular phase observations display an axisymmetric line profile of [O I], red-ward attenuation of the emission of H post 125 days and flattening in the -band marking the onset of dust formation.
@@ 1 Introduction
Massive stars ( 8 , Heger et al., 2003) reach the termination of their evolutionary phase upon exhaustion of their nuclear fuel and undergo core-collapse supernovae (CCSNe). SNe II form a subset of CCSNe that show Balmer features in their spectral sequence (Filippenko, 1997). The progenitors of these SNe are confirmed to be red supergiants (RSGs) through pre-explosion imaging of several SNe II in nearby galaxies (Smartt, 2009; Van Dyk, 2017).
@@
Massive stars are thought to have steady stellar winds driven by radiation pressure (Chiosi & Maeder, 1986). The Galactic RSGs show an observed mass-loss rate in the range of (Beasor & Davies, 2018), due to its dependence on their luminosity and effective temperature. However, early spectral observations of SNe II have shown a short-lived ‘flash’ phase with narrow high-ionization features, indicating a much higher mass loss rate than steady winds (Dessart et al., 2022). Such features were first studied extensively in SN 1998S (Gerardy et al., 2000) as they trace the confined CSM arising from late stages of mass-loss in RSGs (Yaron et al., 2017). A significant fraction of SNe II discovered by Zwicky Transient Facility (ZTF) exhibit such spectroscopic signatures of early interaction during their evolution ( 36%, Bruch et al., 2023).
@@
The effects of a dense CSM driven by high mass loss rates have been known to delay shock breakout (Förster et al., 2018). It occurs when the shock encounters a medium of significantly low optical depth, allowing the radiation to escape in front of the shock (Waxman & Katz, 2017). The breakout is the first electromagnetic signature in CCSNe and happens within a few hours after core-collapse for a typical RSG (Matzner & McKee, 1999). However, if the massive star is engulfed in a compact, dense CSM, the shock breakout leads to an elongated and brighter electromagnetic signature (Ofek et al., 2010; Chevalier & Irwin, 2011), affecting the UV light curves and color evolution (Hiramatsu et al., 2021a; Terreran et al., 2022; Irani et al., 2023). The emergence of high-cadence all-sky surveys (Förster et al., 2018; Subrayan et al., 2023) coupled with prompt follow-up observations has photometrically revealed that fast-rising light curves are primarily reproduced via interaction with CSM. Such an interaction with CSM complicates the classification of SNe II (Singh et al., 2019). SNe II are primarily classified into two main sub-types (Barbon et al., 1979) based on their plateau phase curve decline rates: a Type IIP SN with a slow decline () and a Type IIL SN with a steep linear decline () (Valenti et al., 2016). The primary distinction between Type IIP/L SNe arose from the divide between the outer hydrogen envelope mass (Heger et al., 2003; Eldridge et al., 2018; Hiramatsu et al., 2021b). However, studies by Moriya et al. (2017, 2018); Morozova et al. (2017) have suggested that a higher decline rate during the plateau phase is primarily driven by interaction with CSM, its compactness, and geometry (Andrews et al., 2019; Jacobson-Galán et al., 2022). Chugai (2021) also indicated that higher CSM mass and wind density have been inferred for steeper declining SNe IIL compared to SNe IIP.
@@
Detailed studies of highly interacting SNe II such as SN 2013fs (Yaron et al., 2017), SN 2020pni (Terreran et al., 2022), SN 2020tlf (Jacobson-Galán et al., 2022) etc, have suggested a much higher mass-loss rate for their RSG progenitors, i.e., . Since steady stellar winds cannot propel such a high mass loss rate in an RSG (Puls et al., 2008), the paradigm has shifted towards a more complex scenario in recent years. These mechanisms include eruptive bursts of mass loss (Smith & Owocki, 2006), extensive RSG mass loss through surface instability (Yoon & Cantiello, 2010), wave-driven mass loss during the end stages of nuclear burning (Quataert & Shiode, 2012) and/or turbulent convection driven by dynamical instability (Smith, 2014). The complex mass-loss mechanisms markedly influence the properties of a CCSN through the diverse pathways of forming CSM and its resulting geometrical configuration.
@@
This leads us to the nearest SN II in the decade, SN 2023ixf, on which several studies have been published. SN 2023ixf was discovered by Koichi Itagaki (Itagaki, 2023) at JD 2460084.4 and reported on Transient Name Server (TNS) followed by a prompt classification as a Type II SN (Perley & Gal-Yam, 2023) with visible signatures of flash ionization emission features. Early observational studies detailed the evidence of a shock breakout within a dense CSM (Yamanaka et al., 2023; Jacobson-Galan et al., 2023; Smith et al., 2023; Hosseinzadeh et al., 2023; Bostroem et al., 2023; Teja et al., 2023a; Hiramatsu et al., 2023; Zimmerman et al., 2024; Yang et al., 2024; Li et al., 2024). The evidence of asymmetry in the dense CSM was further discussed by (Smith et al., 2023; Vasylyev et al., 2023; Li et al., 2024). The SN was also detected in X-rays 4 days beyond explosion (Grefenstette et al., 2023; Chandra et al., 2023; Zimmerman et al., 2024) as well as in radio (Matthews et al., 2023) along with tight non-detections in mm-wavelengths (Berger et al., 2023) indicating strong signatures of interaction. Pre-explosion imaging studies (Kilpatrick et al., 2023; Pledger & Shara, 2023; Van Dyk et al., 2023; Neustadt et al., 2024; Jencson et al., 2023; Soraisam et al., 2023; Niu et al., 2023; Qin et al., 2023; Ransome et al., 2024; Xiang et al., 2024) revealed the presence of an RSG progenitor embedded in large amounts of dust with indications of periodic variability. However, the inferred progenitor masses vary significantly, ranging from 8 to 24 . Several other studies have established multi-wavelength limits on progenitor activity (Basu et al., 2023; Dong et al., 2023; Matsunaga et al., 2023; Kong, 2023; Panjkov et al., 2023).
Our work presents the multi-wavelength analysis of SN 2023ixf through hydrodynamic light curve modeling until the early nebular phase and long-term spectroscopic and polarimetric monitoring, highlighting the signatures of asymmetries in SN 2023ixf. The data spanning the first 20 days of our study have already been published in Teja et al. (2023a) discussing the early photometric and spectroscopic analysis. We adopt the explosion epoch as estimated by Teja et al. (2023a), i.e., JD 2460083.315, and the epoch described hereafter in the text is described with respect to this date. We provide an overview of the host galaxy’s properties in Section 2, and summarize the details on observational data acquisition and reduction in Section 3. The photometric properties of SN2023ixf are analyzed and discussed in Section 4, followed by a discussion on its spectroscopic evolution in Section 5. Section 6 delves into the temporal evolution of the polarimetric characteristics. Section 7 describes the hydrodynamical light curve modeling to infer the properties of the progenitor and its multi-faceted CSM. We qualitatively discuss the presence of a confined CSM and the signatures of asphericity in the CSM/ejecta in order to understand the progenitor of SN 2023ixf in Section 8 and summarize our results in Section 9.
@@ 2 Host Galaxy - M101
SN 2023ixf exploded in the outer spiral arm of the face-on spiral Galaxy, M 101 (NGC 5457), located in the constellation Ursa Major at a redshift of z = 0.0008046 obtained from NASA Extragalactic Database (NED111https://ned.ipac.caltech.edu). SN 2023ixf is the 3rd closest CCSN in this millennium after SN 2004dj (Vinkó et al., 2006) and SN 2008bk (Van Dyk et al., 2012). SN 2023ixf lies proximal to one of the giant H II regions in M 101, namely NGC 5461 (see Figure 1), and in the immediate neighborhood of the H II region #1086 (Hodge et al., 1990). Using emission-line diagnostics of the spectra of H II region #1086, (Van Dyk et al., 2023) estimated an oxygen abundance through of 8.43 12 + log[O/H] 8.86, which equates to a metallicity of 0.10 Z 0.20 close to the site of the SN. We adopt a mean distance of 6.82 0.14 Mpc ( = 29.17 0.04 mag, Tikhonov et al., 2015; Riess et al., 2022), a Galactic reddening of = 0.00770.0002 mag (Schlafly & Finkbeiner, 2011), the host reddening of = 0.031 0.011 mag (Lundquist et al., 2023), and a total reddening of = 0.039 0.011 mag for SN 2023ixf.
@@ 3.1 Ultraviolet/Optical/Near-Infrared Photometry
The optical photometric observations of SN 2023ixf were performed using the robotic 0.7-m GROWTH-India telescope (GIT, Kumar et al., 2022) located at the Indian Astronomical Observatory (IAO) in Hanle, India. The observations were carried out in the Sloan Digital Sky Survey (SDSS) filters beginning 2023 May 20 UT. The data were processed with the standard GIT image processing pipeline described in Kumar et al. (2022), and the steps followed are described in Teja et al. (2023a). We also carried out photometric observations in using a 0.36-m Schmidt Cassegrain telescope (Celestron EdgeHD 1400) at the Home observatory in Nayoro, Hokkaido, utilizing a CCD FLI ML1001E camera with an IDAS filter (standard system). The data reduction and aperture photometry were carried out using the software MIRA Pro x64 (Mirametrics, Inc., 2023). The calibration was performed using the stars from APASS catalog (Zacharias et al., 2013). Additionally, optical photometric observations in bands were also carried out using Atik 460 EX Mono CCD mounted on the 0.61-m Vasistha telescope at Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), Sitapur, ICSP, Kolkata. Photometric calibrations were done using Tycho software and the ATLAS catalog (Tonry et al., 2018). -band imaging was also carried out using the 0.51-m telescope at Oku Observatory, Okayama, using the SBIG Camera STXL-6303. The data reduction was performed according to the standard procedure using IRAF.
@@
The Ultraviolet Optical Telescope (UVOT; Roming et al., 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) monitored SN 2023ixf extensively beginning May 21, 2023. We reduced the publicly available data obtained from Swift Archives444Swift Archive Download Portal. Photometry was performed using the UVOT data analysis software in HEASoft, following the procedure described in Teja et al. (2022). Upon using a 5 aperture, a significant part of the nearby H II contaminates the flux, especially in the ultraviolet (UV) bands at later epochs ( 20 d). We used the archival data for the host M 101 obtained on Aug 29, 2006, with OBSID 00035892001, available in the Swift archive, as template images for removing the host contribution. The flux obtained at the SN site in the template images is comparable to the late-phase fluxes in all the swift bands. We employed Swift_host_subtraction555https://github.com/gterreran/Swift_host_subtraction code (Brown et al., 2009, 2014a) to remove the host contribution.
@@ 3.2 Optical Spectroscopy
Low-resolution optical spectroscopic observations of SN 2023ixf were carried out using the Himalayan Faint Object Spectrograph (HFOSC) instrument mounted on the 2-m Himalayan Chandra Telescope (HCT) at IAO (Prabhu, 2014). The HFOSC observations were performed in grisms Gr#7 (3500-7800 Å, R 500) and Gr#8 (5200-9250 Å, R 800). The spectroscopic data were reduced in a standard manner using the packages and tasks in IRAF with the aid of the Python scripts hosted at RedPipe (Singh, 2021). Optical spectroscopic observations were also performed using the Kyoto Okayama Optical Low-dispersion Spectrograph with optical-fiber Integral Field Unit (KOOLS-IFU, Matsubayashi et al., 2019) mounted at the 3.8-m Seimei Telescope (Kurita et al., 2020) located in Okayama Observatory, Kyoto University, Japan. The KOOLS-IFU observations were carried out using VPH-blue (4100-8900 Å, R 500). The data reduction was performed using the Hydra package in IRAF and a reduction software developed for KOOLS-IFU data666http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools. Arc lamps of Hg, Ne, and Xe were used for wavelength calibration. We obtained additional optical spectroscopic data using the TriColor CMOS Camera and Spectrograph (TriCCS) installed on the 3.8-m Seimei telescope having a wavelength coverage of 4000–10500 Å (R 700) with grisms g, r and iz. We used the L. A. Cosmic pipeline (van Dokkum, 2001) to remove cosmic rays. We used arc lamps (Hg, Ne, and Xe) for wavelength calibration, and the flux was calibrated using spectrophotometric standards. All the spectra have been continuum calibrated with respect to photometry and corrected for the redshift of the host galaxy.